# Still More Shades of Null: A Benchmark for Responsible Missing Value Imputation

### v1

@@ 3.3. Evaluation Metrics
Following (Lin and Tsai, 2020; Hasan et al., 2021), we evaluate performance of MVI techniques in two ways: directly using imputation quality metrics and indirectly based on downstream model performance metrics.
3.3.1. Imputation Quality Metrics
Shadbahr et al. (Shadbahr et al., 2023) report that classically used imputation quality metrics are uncorrelated to downstream model performance, and that distributional metrics perform better. To confirm or refute this claim, we use a mix of discrepancy and distributional metrics.
To assess agreement with true values, we compute Root Mean Square Error (RMSE) for numerical and F1 score for categorical attributes. To assess distributional alignment, we compute KL-divergence for both numerical and categorical attributes, for the imputed values as well as for the entire training and test sets. We compute KL-divergence as the Shannon entropy of the probability distributions of the true and predicted values. For categorical attributes, we obtain the probability distributions using the value_counts method with normalization from pandas. For numerical attributes, we first estimate probability density functions (PDFs) using Gaussian kernel density estimation from the scipy package, and then evaluate these estimated PDFs on a set of 1,000 random points to derive the probability distributions. Finally, to assess imputation fairness (Zhang and Long, 2021b, 2022), we compute F1 score difference, RMSE difference, and KL divergence difference between privileged and disadvantaged groups.
3.3.2. Model Performance Metrics
To evaluate model performance we utilize Virny (Herasymuk et al., 2024), a comprehensive software library that computes fairness, stability, and accuracy metrics for a set of initialized models, on several single-attribute and intersectional subgroups of interest. Specifically, to assess the impact of MVI techniques on model correctness, we report the F1 score because it is a more reliable metric than accuracy for imbalanced data. For evaluating model stability, we report average Label Stability (Darling and Stracuzzi, 2018; Khan et al., 2023) over the full test set (closely related to the self-consistency metric from Cooper et al. (Cooper et al., 2023)), computed per-sample for binary classification Label Stability = $\frac{|B_{+}-B_{-}|}{B}$, where $B_{+}$ is the number of times the sample is classified into the positive class and $B_{-}$ is the number of times the sample is classified into the negative class, and $B=B_{+}+B_{-}$ models are trained by bootstrapping over the train set. We set $B=50$ in all of our experiments.
Lastly, we compute standard fairness metrics based on group-specific error rates, namely True Positive Rate Difference (TPRD), True Negative Rate Difference (TNRD), Selection Rate Difference (SRD), and Disparate Impact (DI). (Note that DI computes the ratio of selection rates, but we refer to it as DI as is standard in the literature (Feldman et al., 2015).)
@@ 3.3.2. Model Performance Metrics
3.4. Shades-of-Null Benchmark Architecture
Figure 1 summarizes the architecture of the Shades-of-Null benchmark. Our goal when developing this benchmark was to support a large-scale evaluation of the impact of different missingness scenarios (Section 3.1) and missing value imputation techniques (Section 3.2) on different aspects of performance of machine learning models (Section 3.3). This benchmark is novel in several ways. First, it handles both single-mechanism and mixed missingness scenarios, including missingness shift. Second, it supports flexible evaluation of imputation quality and model performance — including accuracy, fairness and stability — for multiple sensitive attributes and protected groups, as well as their intersections.
The central component of the architecture in Figure 1 is the benchmark controller that executes the specified evaluation scenarios and applies error injectors to input datasets, accordingly. It then imputes missing values and preprocesses the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features. The controller then conducts ML model hyperparameter tuning. Finally, the evaluation module assesses MVI and ML model performance, both overall and for specific subgroups. To facilitate comprehensive model profiling, the evaluation module utilizes the Virny software library (Herasymuk et al., 2024), measuring correctness, fairness and stability. We store experimental results in MongoDB.
Figure 1. Shades-of-Null benchmark architecture
@@ 3.4. Shades-of-Null Benchmark Architecture
folk finance 15,000 sex, race
@@
Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments. Firstly, it separates missing value imputation and model evaluation stages. This enables the preservation of imputation results, including imputed datasets stored in a file system and imputation performance metrics stored in a database, facilitating their reuse during the subsequent model evaluation stage. Consequently, redundant application of the same MVI on identical datasets with identical seeds for the pipeline is avoided when assessing different models. Secondly, the experimental pipeline within the benchmark is tailored to evaluate MVI and ML models simultaneously on multiple test sets, with varying error rates and missingness types. As both MVI and ML models only perform inference on test sets, the running time for executing a pipeline with one training set and one test set is close to that of a pipeline with one training set and multiple test sets.
@@ 3.5. Datasets and Tasks
A summary of the datasets used in this study is given in Table 2.
@@
german [^3] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) contains records of creditworthiness assessments, classifying individuals as high or low credit risks. It contains information on 1,000 individuals characterized by 20 attributes (13 categorical, 7 numerical). Here, sex and age are the sensitive attributes, with “female” and ¡25 as the disadvantaged groups.
@@
folk (Ding et al., 2021) was derived from US Census surveys conducted in 2014-2018, and is associated with several tasks. We selected ACS Income, a binary classification task to predict whether a person’s annual income is above \$50,000. We use 2018 data from Georgia, with 15,000 tuples and 10 attributes (8 categorical, 2 numerical). Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
@@
law-school [^4] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains law school admissions records of 20,798 applicants, characterized by 12 attributes (6 categorical, 6 numerical). The task is to predict whether a candidate would pass the bar exam. Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
@@
bank [^5] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical) and a binary target that represents whether the individual signed up for the product. Here, age is the sensitive attribute, with ¡25 and ¿60 as the disadvantaged group.
@@
heart [^6] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, wiht information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), and a target variable that denotes the presence of a heart disease. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
4. Single-Mechanism Missingness
In our first set of experiments, we compare the performance of MVI techniques in single-missingness scenarios (S1-S3 in Table 1), with the same fraction of errors (30%) in train and test.
@@ 4. Single-Mechanism Missingness
For correctness, we report results for F1, and defer additional results to Appendix C.2. For fairness, we report metrics based on an intersectional group (e.g., sex&race) for datasets with two sensitive attributes, and a single group (e.g., age) for datasets with one sensitive attribute. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the train set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
We evaluate 5 ML types: Decision Tree (DT), Logistic Regression (LR), Gradient Boosted Trees (LGBM), Random Forest (RF), and a Neural Network (historically called the Multi-Layer Perceptron or MLP). Different models may perform best on different datasets. In Figures 34 and 5, we report on the best-performing models (based on their F1 score) for each dataset, and make it explicit that we are comparing across different model types. Additionally, we compare performance against a model that is trained on the clean train set using the same model architecture.
4.1. Correctness of the Predictive Model
As shown in Figure 3, the efficacy of MVI techniques, and the downstream effect on the predictive performance of the model, according to F1, depends significantly on the characteristics of the dataset and the missingness mechanism.
All the methods we evaluated are competitive for all missingness mechanisms on heart and law-school. boostclean, which uses multiple imputation (MI), is competitive only on small datasets (diabetes and german), and only under MCAR. It performs particularly poorly on folk, with as much as a 0.8 drop in F1 compared to other methods. We discuss this unexpected performance of MI further in Section 7. In line with conventional wisdom, deletion significantly worsens predictive performance for small datasets. For diabetes (the smallest dataset), F1 drops by 0.06 for MCAR, 0.08 for MAR and 0.1 for MNAR, compared to ML model trained on clean data. Under MCAR, miss-forest (ML-based) and auto-ml (DL-based) are the best-performing techniques. Under MAR and MNAR there is no clear best-performing technique. For example, under MAR datawig ( DL-based) shows good performance, but simple median-mode and median-dummy imputers are also competitive. This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (on a variety of missingness mechanisms) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
@@ 4.1. Correctness of the Predictive Model
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
4.2. Fairness of the Predictive Model
Next, we evaluate the fairness of the downstream ML model for different MVI techniques. We present TPRD in Figure 4, with results for other fairness metrics in Appendix C.2. The effect of MVI on fairness (error-disparity) is highly dataset specific, and it is strongly correlated with fairness of the model trained on clean data (no missingness), corroborating the findings of Guha et al. (Guha et al., 2024).
On heart and german, we observe negligible disparity for the model trained on clean data, and MVI techniques do not further exacerbate it. On folk and bank, we see large disparities (close to 20% and 10%, respectively, for model trained on clean data). Interestingly all MVI techniques preserve or improve fairness (reduce disparity) in all missingness settings, with boostclean showing lowest disparity. On diabetes and law-school (which, for the model trained on clean data, show disparities close to 10% and 8%, respectively), MVI techniques such as boostclean, clustering and deletion can worsen fairness, with the largest effect under MNAR.
Wang and Singh (2021) posit that models will exhibit more unfairness under MAR and MNAR compared to MCAR, but we only find weak empirical evidence towards this, even for techniques like deletion. A nuance here is in the way we set up the missingness scenarios, described in detail in Section 3.1. We took a broad socio-technical lens and designed non-response scenarios that are realistic — including MAR scenarios where people from disadvantaged groups withhold information that could hurt their chances of getting the desired outcome. Hence, dropping these rows might in fact improve “fairness” by removing feature values that contribute negatively to outcomes for the disadvantaged group. With this intuition, deletion can in fact improve fairness under MAR and MNAR, as observed on bank and diabetes.
@@ 4.2. Fairness of the Predictive Model
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are ordered in increasing order by size. Blue line shows performance of the model trained on clean data.
@@ 4.3. Stability of the Predictive Model
Next, we compare the stability of the downstream model trained using different MVI techniques. Results in Figure 5 show that, as expected, stability depends most strongly on dataset characteristics (mainly size), then on imputation strategy.
All imputers preserve or marginally improve stability on bank and heart (largerst datasets, with 40k and 70k tuples, respectively), for all missingness mechanisms. Even deletion does not worsen stability in this case since there is sufficient data even after dropping rows with nulls, and clustering marginally improves stability over the model trained on clean data, for all missingness types.
On folk and law-school (medium-sized datasets, with 15k and 20k tuples, respectively), most MVI techniques are competitive, except for boostclean, which shows a 5% drop in label stability compared to the model trained on clean data. boostclean does joint cleaning and model training, and is expected to be sensitive to small changes in the training data, as validated by our findings.
On german and diabetes (about 1K samples each) deletion, clustering and boostclean are the worst performing methods on all missingness mechanisms, with label stability 8-10% lower compared to the model trained on clean data. miss-forest and auto-ml are the best performing methods, and match the stability of the model trained on clean data for all missingness types.

### v2

@@ 3.3. Evaluation Metrics
Following (Lin and Tsai, 2020; Hasan et al., 2021), we evaluate MVI techniques in two ways: directly using imputation quality metrics and indirectly based on downstream model performance.
3.3.1. Imputation Quality
Shadbahr et al. (Shadbahr et al., 2023) report that distributional metrics capture downstream model performance better than classically-used discrepancy metrics. To confirm or refute this claim, we use a mix of both. To assess agreement with true values, we compute Root Mean Square Error (RMSE) for numerical features and F1 score for categorical features. To assess distributional alignment, we compute KL-divergence (i.e., the Shannon entropy) between the true and the predicted feature distributions, for both numerical and categorical features, measured for the imputed columns only as well as for the full dataset. For categorical features, we obtain the probability distributions using the value_counts method with normalization from pandas. For numerical features, we use Gaussian kernel density estimation from scipy, with 1000 samples. Finally, to assess imputation fairness (Zhang and Long, 2021b, 2022), we compute F1 score difference, RMSE difference, and KL divergence difference between privileged (priv) and disadvantaged (dis) groups.
3.3.2. Model Performance
To assess the impact of MVI techniques on model correctness, we report the F1 score because it is a more reliable metric than accuracy for imbalanced data.
For evaluating model stability, we report average Label Stability (Darling and Stracuzzi, 2018; Khan et al., 2023) over the full test set (closely related to the self-consistency metric from  Cooper et al. (2023)), computed per-sample for binary classification as Label Stability = $\frac{|B_{+}-B_{-}|}{B}$, where $B_{+}$ is the number of times the sample is classified into the positive class and $B_{-}$ is the number of times the sample is classified into the negative class, and $B=B_{+}+B_{-}$ models are trained by bootstrapping over the train set. We set $B=50$ in all our experiments.
Lastly, we report model fairness based on group-specific error rates, namely True Positive Rate Difference (TPRD), True Negative Rate Difference (TNRD), Selection Rate Difference (SRD), and Disparate Impact (DI). (Note that DI computes the ratio of selection rates, but we refer to it as DI as is standard in the literature (Feldman et al., 2015).)
@@ 3.3.2. Model Performance
3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. Our goal when developing this evaluation suite was to support the comprehensive evaluation of different MVI techniques (Section 3.2), under different missingness scenarios (Section 3.1), according to different aspects of model performance (Section 3.3). Shades-of-Null is novel in several ways. First, it is flexible and supports a variety of missingness scenarios, namely, single-mechanism, multi-mechanism and missingness shift. Second, it supports holistic evaluation of both imputation quality (including imputation fairness) and model performance — including accuracy, stability and fairness.
The central component of the architecture in Figure 1 is the benchmark controller that executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values according to user-specified MVI technique(s) and pre-processes the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features. The controller then executes ML model training, including hyperparameter tuning. Finally, the evaluation module assesses imputation quality and model performance. For comprehensive model profiling, the evaluation module uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability and fairness metrics for a set of initialized models, for multiple sensitive attributes and their intersections. We store experimental results in MongoDB.
Figure 1. Shades-of-Null architecture
@@ 3.4. Shades-of-Null Architecture
folk-income finance 15,000 sex, race
@@
folk-employment hiring 302,640 sex, race
Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments. Firstly, it decouples missing value imputation and model training. This enables the preservation of imputation results, including imputed datasets stored in a file system, facilitating their reuse during subsequent model training and evaluation stages. Secondly, the experimental pipeline within Shades-of-Null supports evaluation on multiple test sets simultaneously (e.g., with different rates or types of missingness), to speed up evaluation. For example, the running time of executing a pipeline with one training set and one test set is close to that of a pipeline with one train set and multiple test sets.
@@ 3.5. Datasets and Tasks
As noted in Section 1, in this work we are interested in missingness that is socially salient. With this in mind, we selected $7$ datasets from a variety of social decision-making contexts including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2. For each dataset, there is a binary classification task, where being assigned a positive label corresponds to the attainment of a desirable social good (such as education, employment, or access to healthcare). We selected these datasets to have good coverage of (i) social domains and (ii) dataset sizes and other characteristics.
@@
german [^4] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) is a popular fairness dataset that contains records of creditworthiness assessments, classifying individuals as high or low credit risks. It contains information on 1,000 individuals characterized by 20 attributes (13 categorical, 7 numerical), including credit history, occupation and housing information. Here, sex and age are the sensitive attributes, with “female” and “age25absent25$\leq$ 25≤ 25” as the disadvantaged groups.
@@
Folktables (Ding et al., 2021) is another popular fairness dataset derived from US Census data from all 50 states between 2014-2018. The dataset has several associated tasks, of which we selected two: (i) ACSIncome (folk-income) is a binary classification task to predict whether an individual’s annual income is above \$50,000, from 10 features (8 categorical, 2 numerical) including educational attainment, work hours per week, marital status, and occupation. We use data from Georgia from 2018, subsampled to 15k rows. (ii) ACSEmployment (folk-employment) is a binary classification task to predict whether an individual is employed, from 17 features (16 categorical, 1 numerical) including educational attainment, employment status of parent, military status, and nativity. We use data from California from 2018, with 302,640 rows (we do not subsample). In both tasks, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
@@
law-school [^5] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains admissions records of 20,798 applicants, characterized by 12 attributes (6 categorical, 6 numerical), including LSAT scores and college GPAs. The task is to predict whether a candidate would pass the bar exam. Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
@@
bank [^6] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit. Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
@@
heart [^7] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, with information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), including age, height, weight, blood pressure, and a binary target indicating whether the patient has heart disease. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
3.6. Model Types
We evaluate predictive performance of 6 ML models: (i) decision tree (dt__\__clf) with a tuned maximum tree depth, minimum samples at a leaf node, number of features used to decide the best split, and criteria to measure the quality of a split; (ii) logistic regression (lr__\__clf) with tuned regularization penalty, regularization strength, and optimization algorithm; (iii) gradient boosted trees (lgbm__\__clf) with tuned number of boosted trees, maximum tree depth, maximum tree leaves, and minimum number of samples in a leaf; (iv) random forest (rf__\__clf) with a tuned number of trees, maximum tree depth, minimum samples required to split a node, and minimum samples at a leaf node (v) neural network, historically called the multi-layer perceptron (mlp__\__clf) with two hidden layers, each with 100 neurons, and a tuned activation function, optimization algorithm, and learning rate; (vi) a deep table-learning method called GANDALF (Joseph and Raj, 2022) (gandalf__\__clf) with a tuned learning rate, number of layers in the feature abstraction layer, dropout rate for the feature abstraction layer, and initial percentage of features to be selected in each Gated Feature Learning Unit (GFLU) stage.
4. Single and Multi-Mechanism Missingness
To simulate single-mechanism missingness (S1-S3 in Table 1) we inject 30% of each training and test sets with nulls, according to the missingness scenarios described in Section 3.1. For multi-mechanism or mixed missingness, when MCAR, MAR and MNAR co-exist (S10 in Table 1), we inject 10% of nulls for each of the three mechanism into both training and test sets, for a total of 30% nulls.
@@ 4. Single and Multi-Mechanism Missingness
To evaluate model correctness, we report results for F1, see Appendix C.1 for accuracy results. For fairness, we use binary group definitions throughout. For datasets with two sensitive attributes, we define the doubly-disadvantaged group as disadvantaged (dis) and everyone else as privileged (priv). For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men collectively form the priv group. We report results for TPRD, and defer results for other fairness metrics to Appendix C.1. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the training set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
Different models are the best-performing on different datasets. In Figures 34 and 5, we report on the best-performing models (according to F1) for each dataset, and compare performance against a model trained on the clean training set using the same model architecture (performance shown as a blue line in the plots).
@@
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are in increasing order by size. Blue line shows performance of the model trained on clean data.
4.1. Correctness of the Predictive Model
Figure 3 shows the F1 of models trained with different MVI techniques. We find interesting trends in MVI performance based on characteristics of the dataset and missingness type. All techniques are competitive for all missingness mechanisms, including mixed missingness, on heart and law-school. boostclean, which uses multiple imputation (MI), is otherwise only competitive on small datasets (diabetes and german), and only under MCAR and mixed missingness on german. None of the MVI techniques are able to match the F1 of the model trained on clean data on folk-employment, and this effect is strongest under MNAR (notably, stronger than under mixed missingness). boostclean shows particularly poor performance on folk-income, with a 0.08 decrease in F1 compared to other methods, for all missingness types. We discuss this unexpected performance of MI further in Section 6.
auto-ml, datawig and miss-forest are generally the best performing MVI techniques, but simpler statistical techniques (e.g., median-mode under MAR) are also competitive. This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
In line with conventional wisdom (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we find that deletion worsens predictive performance. This effect is strongest for small datasets like diabetes, with F1 decreasing as much as 0.1 under MNAR, compared to the model trained on clean data.
Note that the F1 score on the bank dataset is extremely low (0.32). This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 8 in Appendix B). Further, from Figure 18 in Appendix C.1, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15.
@@ 4.1. Correctness of the Predictive Model
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are in increasing order by size. Blue line shows TPRD of the model trained on clean data.
4.2. Fairness of the Predictive Model
Figure 4 shows the effect of MVI on fairness, according to TPRD.  Wang and Singh (2021) posit that models will exhibit more unfairness under MAR and MNAR compared to MCAR, but we only find weak empirical evidence towards this, even for deletion. A nuance here is that we designed missingness scenarios, described in Section 3.1, to be realistic — including MAR scenarios where people from disadvantaged groups withhold information that could hurt their chances of getting the desired outcome. Hence, dropping these rows can in fact improve fairness under MAR and MNAR, as observed on bank.
In contrast to Wang and Singh (2021), we find that the effect of MVI on fairness is strongly correlated with fairness of the model trained on clean data, corroborating the findings of  Guha et al. (2024). Fairness depends on two things: dataset characteristics and model type. All MVI techniques except for boostclean have the same model type as the model trained on clean data (because they are impute-then-classify approaches) and generally preserve fairness of that model, under all missingness mechanisms. Notably, this is agnostic to whether the TPRD of the model trained on clean data is low (close to 0.01 on heart and folk-employment, and 0 on german) or high (close to -0.1 on folk-income and 0.2 on bank).
On the other hand, boostclean, is a joint data cleaning and model training approach and thereby constitutes it own model type, and shows fairness trends that deviate from the model trained on clean data. boostclean significantly improves fairness on
folk-income (TPRD close to 0, compared to -0.1 for the clean model) and bank (TPRD close to 0.1, compared to 0.2 for the clean model), but marginally worsens fairness on law-school (TPRD -0.1 compared to -0.08 for the clean model) and heart (TPRD -0.02 compared to 0.01 for the clean model), for all missingness types.
@@ 4.3. Stability of the Predictive Model
Figure 5 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german. boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank and all missingness types: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
(a) F1 (imputation) vs F1 (model)
(b) RMSE (imputation) vs F1 (model)
(c) KL divergence (imputation) vs F1 (model)
Figure 6. Imputation quality vs. model performance: imputation correctness (F1, RMSE and KL divergence) may not be indicative of model correctness (F1).